OpenStudy (anonymous):

NEED HELP PLEASE...having problem solving. given the function f(x)=3x^3-9x^2-2, find the coordinates of all the critical points of the function.

OpenStudy (anonymous):

first differentiate

OpenStudy (anonymous):

i did that and got 9x^2 - 18 = 0, but this where i'm stuck

OpenStudy (anonymous):

i think the derivative of thefunction is not 9x^2 - 18 = 0, but 9x^2-18x=0 x^2-2x=0 x(x-2)=0 x=2 and x=0

OpenStudy (anonymous):

then plug the values of x to the function f(0) = -2 f(2)= 3(2)^3 -9(2)^2 -2 = 24-36-2 = -14 so the critical points is (0,-2) and (2,-14)

OpenStudy (anonymous):

oh ok. i got that! thanx alot